OpenStudy (anonymous):

A square and a rectangle have the same perimeter of 108 cm. The length of the rectangle is 6 cm greater than the average of all its sides. Which of the two figures has a greater area and by how many cm2?

OpenStudy (anonymous):

Average of all sides of rectangle = (2L + 2W) / 4 = (L+W)/2 L = (L+W)/2 + 6 => 2L - 12 = L + W We also know L + W = 108/2 => L + W = 54 So, 2L - 12 = 54 => 2L = 66 => L = 33 => W = 21 Area = 33 * 21 = 693 sq cm Area of square = 27*27 (since each side is 108/4) = 729 sq cm.

OpenStudy (anonymous):

Thanks so much!!! This helped me a lot!!